OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1=-12 and a27=66.

OpenStudy (campbell_st):

you need to find the common difference d using the 27th term \[T _{27} = -12 + (27 -1)d\] or \[66 = -12 + 26d\] solve for d then evaluate the formula with n = 42

OpenStudy (anonymous):

\[a(n) = a(1) + (n - 1)d\] \[a(1) = -12\]a(n) = -12 + (n - 1)d \[a(27) = 66\]66 = -12 + (27 - 1)d \[66 = -12 + 26d\] \[26d = 66 + 12\] \[26d = 78\] \[d = 78/26\] \[d = 3\] \[a(n) = -12 + 3(n - 1)\] \[a(42) = -12 + 3(42 - 1) = -12 + 3(41) = -12 + 123 = 111\]